# Entropy-Gated Branching for Efficient Test-Time Reasoning

### v3

@@ Abstract
Test-time compute methods can significantly improve the reasoning capabilities and problem-solving accuracy of large language models (LLMs). However, these approaches require substantially more computational resources, with most compute wasted on exploring low-diversity branches where the model already exhibits high confidence. We observe that a small subset of uncertain reasoning steps has a disproportionately large impact on final prediction accuracy, and branching at these critical junctures tends to yield more diverse and higher-quality candidate reasoning steps. We propose Entropy-Gated Branching (EGB), which branches only at high-uncertainty steps and prunes expansions with a lightweight verifier. On mathematical and financial reasoning benchmarks, EGB improves accuracy by 22.6% over standard inference while operating 31%-75% faster across math benchmarks than test-time beam search with higher performance. Our results show that dynamic resource allocation during inference can substantially improve both efficiency and effectiveness, offering a more scalable pathway to enhanced LLM reasoning capabilities.
@@
Xianzhi Li1, Ethan Callanan1, Abdellah Ghassel1, Xiaodan Zhu1, 1Queen’s University {li.xianzhi, e.callanan, abdellah.ghassel, xiaodan.zhu}@queensu.ca
@@ 4 Experiments
We implement our method on two families of models, Qwen3 Yang et al. (2025) and Llama3 Touvron et al. (2023), at 3 sizes each. For Qwen3 we use Qwen3-1.7B, Qwen3-4B, and Qwen3-8B, all in "non-thinking" mode Yang et al. (2025). For Llama3 we use Llama-3.2-1B-instruct, Llama-3.2-3B -instruct, and Llama-3.1-8B-instruct. Unless otherwise noted, we adopt the recommended decoding settings: for Qwen3, temperature 0.7, top-p 0.8, top-k 20, and min-p 0; for Llama3, temperature 0.6 and top-p 0.9.
@@
To provide reliable feedback, we use Qwen2.5-Math-PRM-7B, which is the most popular PRM model on HuggingFace up to date. This model is specifically trained to assess reasoning quality by assigning a numerical score that reflects the correctness of the inference. Unlike the instruct model, which relies on a qualitative analysis to provide a branch selection, this model directly provides a probabilistic evaluation, where a designated probability serves as a quantitative score for ranking the branches. Higher scores indicate stronger logical consistency and correctness.
@@
To balance performance with computational overhead, we set the entropy thresholds $\tau=1.5$ for all Qwen models, and $\tau=0.8$ for the Llama models since they exhibit less uncertainty. For results reported in Table 1, we use a beam size $K$ of 4, a beam width $W$ of 4 for all datasets, and a sample budget of 16 for Self-consistency. Detailed analyses of the threshold values and branch numbers can be found in Section A.2.
@@ 5.1 Main results
EGB consistently outperforms standard decoding on all domains, but there are nuanced patterns between models and domains. For smaller models (1-2B parameters), EGB yields comparable improvements across both domains, suggesting that when base model capabilities are limited, both financial and mathematical reasoning benefit equally from increased exploration. However, as model scale increases, a clear trend emerges: financial reasoning tasks maintain robust gains from EGB while general mathematical problems show diminishing returns. At the largest scales tested (8B parameters), financial tasks benefit at approximately twice the rate of general mathematical problems.
This scale-dependent pattern likely reflects fundamental differences in how uncertainty manifests across these reasoning domains. Financial reasoning problems typically require integrative processing that combines numerical computation, contextual interpretation, and domain-specific procedural knowledge. This multi-faceted nature creates what we term a distributed uncertainty landscape, where multiple decision points throughout the reasoning chain contribute meaningfully to solution quality. In such scenarios, EGB’s exploration strategy can effectively navigate the complex interaction between different reasoning components.
@@ 6 Conclusion
While EGB demonstrates substantial improvements across multiple benchmarks, several considerations warrant acknowledgment. First, the method’s effectiveness depends on the quality of the external process reward model used for branch evaluation. Variations in PRM quality may influence results in different deployment contexts. Second, the entropy threshold requires model-specific tuning, with optimal values differing between general mathematical problems and complex financial reasoning tasks. Automated threshold selection mechanisms could enhance generalizability. Finally, our evaluation focuses on mathematical and financial reasoning; further investigation is needed to assess EGB’s effectiveness on other reasoning domains, such as commonsense reasoning or open-ended problem-solving.
@@ A.1 Algorithm
Final answer correctness was determined using a powerful external language model, Gemini 2.5 Flash, as an impartial judge Gemini Team et al. (2024). This approach circumvents the limitations of rigid evaluation methods like exact string matching or regex-based parsers, which are often brittle and can penalize semantically correct answers due to minor stylistic or formatting variations. For each problem, the LLM judge was provided with the original question, the model’s complete generated response, and the ground-truth solution, and was then prompted to assess whether the final answer was correct. This method ensures a robust and fair assessment by focusing on the semantic correctness of the reasoning outcome rather than superficial syntactic differences, providing a more reliable measure of the model’s true problem-solving capabilities.
@@ A.2 Ablation Studies
Threshold Sensitivity
We explore how the entropy threshold affects model performance on different scales. By analyzing the effect of altering the entropy threshold $H_{\text{thresh}}$ from 1.0 to 4.0 on a subset of 100 questions from MATH-500 dataset, Figure 5 shows that $H_{\text{thresh}}=2$ is, on average, the best setting for the Llama models, where performance remains relatively stable within the range of 1.5-3.5. This optimal range highlights that EGB’s effectiveness depends critically on maintaining sufficient branching opportunities during reasoning while avoiding excessive computational overhead. At very low thresholds ($H_{\text{thresh}}<1.5$), the model branches too frequently, often at decision points where the uncertainty is minimal and exploration provides little benefit. This aggressive branching not only increases computational costs but can also introduce noise by exploring fundamentally similar reasoning paths that differ only in minor details or presentation.
Figure 5: Entropy threshold sensitivity analysis using Llama models on MATH-500 subset (20% questions)

### v4

@@ Abstract
Test-time compute methods can significantly improve the reasoning capabilities and problem-solving accuracy of large language models (LLMs). However, these approaches require substantially more computational resources, with most compute wasted on exploring low-diversity branches where the model already exhibits high confidence. We observe that a small subset of uncertain reasoning steps has a disproportionately large impact on final prediction accuracy, and branching at these critical junctures tends to yield more diverse and higher-quality candidate reasoning steps. We propose Entropy-Gated Branching (EGB), which branches only at high-uncertainty steps and prunes expansions with a lightweight verifier. On mathematical and financial reasoning benchmarks, EGB improves accuracy by 22.6% over standard inference while operating 31%-75% faster across math benchmarks than test-time beam search with higher performance. Our results show that dynamic resource allocation during inference can substantially improve both efficiency and effectiveness, offering a more scalable pathway to enhanced LLM reasoning capabilities. We release our code and tools here [^1] https://github.com/JXL884/entropy_gated_branching
@@
Xianzhi Li, Ethan Callanan, Abdellah Ghassel, Xiaodan Zhu Department of Electrical and Computer Engineering & Ingenuity Labs Research Institute Queen’s University {li.xianzhi, e.callanan, abdellah.ghassel, xiaodan.zhu}@queensu.ca
@@ 4 Experiments
We implement our method on two families of models, Qwen3 Yang et al. (2025) and Llama3 Touvron et al. (2023), at 3 sizes each. For Qwen3 we use Qwen3-1.7B, Qwen3-4B, and Qwen3-8B, all in "non-thinking" mode Yang et al. (2025). For Llama3 we use Llama-3.2-1B-instruct, Llama-3.2-3B -instruct, and Llama-3.1-8B-instruct. Unless otherwise noted, we adopt the recommended sampler settings. In addition, we explore thinking-mode decoding for Qwen3 in Appendix A.2. This setting produces substantially longer reasoning traces and different efficiency–accuracy trade-offs.
@@
To provide reliable feedback, we use Qwen2.5-Math-PRM-7B, which is the most popular PRM model on HuggingFace up to date. This model is specifically trained to assess reasoning quality by assigning a numerical score that reflects the correctness of the inference. Unlike the instruct model, which relies on a qualitative analysis to provide a branch selection, this model directly provides a probabilistic evaluation, where a designated probability serves as a quantitative score for ranking the branches. Higher scores indicate stronger logical consistency and correctness. To assess robustness to verifier choice, we additionally evaluate EGB with an alternative PRM, RLHFlow/Llama3.1-8B-PRM-Deepseek-Data, in Appendix A.2.
@@
For results reported in Table 1, we use a beam size $K$ of 4, a beam width $W$ of 4 for all datasets, and a sample budget of 16 for Self-consistency. Detailed analyses of the threshold values and branch numbers can be found in Section A.2.
@@ 5.1 Main results
EGB consistently outperforms standard decoding on all domains, but there are nuanced patterns between models and domains. For smaller models (1-2B parameters), EGB yields comparable improvements across both domains, suggesting that when base model capabilities are limited, both financial and mathematical reasoning benefit equally from increased exploration. However, as model scale increases, financial reasoning tasks maintain robust gains from EGB while general mathematical problems show diminishing returns. At the largest scales tested (8B parameters), financial tasks benefit at approximately twice the rate of general mathematical problems.
This scale-dependent pattern likely reflects fundamental differences in how uncertainty manifests across these reasoning domains. Financial reasoning problems typically require integrative processing that combines numerical computation, contextual interpretation, and domain-specific procedural knowledge. In such scenarios, EGB’s exploration strategy can effectively navigate the complex interaction between different reasoning components.
@@ 6 Conclusion
While EGB demonstrates consistent improvements across mathematical and financial reasoning benchmarks, a few considerations remain. First, although our threshold sensitivity analysis reveals broad "safe zones" where performance is stable, automated threshold selection mechanisms could further enhance ease of deployment. Second, our evaluation focuses on domains where process reward models are well-established; extending EGB to domains with less mature verifiers (e.g., open-ended generation or commonsense reasoning) remains future work. Finally, combining EGB with thinking-mode decoding incurs higher computational costs due to longer reasoning traces, requiring practitioners to balance accuracy gains against latency constraints in resource-sensitive applications.
Acknowledgement
This research was funded in part by the Faculty Research Awards of J.P. Morgan AI Research. The authors are solely responsible for the contents of the paper and the opinions expressed in this publication do not reflect those of the funding agencies. We also thank Mathieu Sibue, Antony Papadimitriou, Zhiqiang Ma, and Xiaomo Liu for helpful discussions and valuable feedback that contributed to this work.
@@ A.1 Algorithm
Final answer correctness was determined using a powerful external language model, Gemini 2.5 Flash, as an impartial judge Gemini Team and others (2024). This approach circumvents the limitations of rigid evaluation methods like exact string matching or regex-based parsers, which are often brittle and can penalize semantically correct answers due to minor stylistic or formatting variations. For each problem, the LLM judge was provided with the original question, the model’s complete generated response, and the ground-truth solution, and was then prompted to assess whether the final answer was correct. This method ensures a robust and fair assessment by focusing on the semantic correctness of the reasoning outcome rather than superficial syntactic differences, providing a more reliable measure of the model’s true problem-solving capabilities.
@@ A.2 Ablation Studies
Base Model PRM Method CFA-L1 CFA-L2 MATH-500 GSM8K Avg
Llama-1B Llama3.1-8B-PRM Beam Search 41.06 32.95 39.60 57.39 42.75
EGB 42.22 33.66 39.40 62.40 44.42
Qwen2.5-Math-PRM Beam Search 43.83 31.82 45.80 63.00 46.11
EGB 42.28 34.09 44.20 61.92 45.62
Qwen3-1.7B Llama3.1-8B-PRM Beam Search 58.91 47.16 71.19 86.03 65.82
EGB 60.80 50.98 75.06 87.81 68.66
Qwen2.5-Math-PRM Beam Search 58.72 50.57 73.80 86.58 67.42
EGB 61.50 51.70 77.40 89.46 70.02
Table 2: Sensitivity to PRM choice. Accuracy of Beam Search and EGB under two process reward models: Qwen2.5-Math-PRM-7B (default in main experiments) and an alternative verifier RLHFlow/Llama3.1-8B-PRM-Deepseek-Data (shown as Llama3.1-8B-PRM). Bold marks the better method within each base-model/PRM setting for each benchmark.
PRM Quality Sensitivity and Domain Generality
We use Qwen2.5-Math-PRM-7B as the process reward model for candidate scoring and ranking throughout our main experiments. In this study, we assess the robustness of EGB to the choice of PRM by replacing it with an alternative verifier, RLHFlow/Llama3.1-8B-PRM-Deepseek-Data, which is trained on different data and may exhibit different calibration properties. We evaluate two base models (Llama-1B and Qwen3-1.7B) on four benchmarks (CFA-L1, CFA-L2, MATH-500, GSM8K), while keeping the same decoding/branching configuration as in the main results (beam size $K=4$, beam width $W=4$). In all settings, the PRM is used only to score and rank candidates, consistent with Algorithm 1.
Table 2 reports accuracy for Beam Search with PRM feedback and EGB under both PRMs. We observe that swapping the PRM can affect absolute performance, particularly on math benchmarks like MATH-500, GSM8K, indicating that verifier quality/calibration impacts downstream selection. Importantly, EGB remains effective under both PRMs. For Qwen3-1.7B, the gain of EGB over Beam Search is consistent across PRMs (average +2.84pp with Llama3.1-8B-PRM and +2.60pp with Qwen2.5-Math-PRM across the four benchmarks). For Llama-1B, EGB improves over Beam Search with Llama3.1-8B-PRM (average +1.67pp), and is comparable under Qwen2.5-Math-PRM (average -0.49pp), with improvements on CFA-L2 but decreases on some math benchmarks. Overall, these results suggest that PRM choice primarily affects absolute accuracy, while entropy-gated branching provides a largely complementary benefit to verifier strength.
Model Task $\tau{=}0$ $0.5$ $1.0$ $1.5$ $2.0$ $2.5$ $3.0$ $4.0$ $\tau{=}\infty$
Qwen3-1.7B CFA-L2 50.57 49.43 47.16 52.27 50.00 48.30 44.89 44.59 44.89
MATH-500 73.80 64.33 74.94 75.00 74.80 77.20 77.60 76.40 74.23
Qwen3-4B CFA-L2 52.27 56.25 56.82 56.25 55.68 57.95 56.82 56.09 56.16
MATH-500 81.20 73.06 80.97 82.58 83.88 84.51 83.40 81.89 78.60
Qwen3-8B CFA-L2 57.39 54.55 57.95 61.36 61.36 58.52 58.52 61.93 59.66
MATH-500 82.40 69.00 79.81 82.74 84.75 84.79 84.91 83.47 78.40
Llama-1B CFA-L2 31.82 27.84 26.70 33.52 31.82 32.95 35.80 36.36 31.82
MATH-500 45.80 42.60 42.20 40.00 39.40 37.58 38.81 36.50 42.80
Llama-3B CFA-L2 38.64 42.05 44.32 38.64 39.77 39.77 42.84 44.32 40.89
MATH-500 55.80 53.22 54.20 54.80 56.13 57.98 59.56 55.87 56.39
Llama-8B CFA-L2 43.75 44.32 42.05 47.16 52.50 55.11 51.70 53.98 50.89
MATH-500 60.20 62.91 61.18 61.41 61.76 66.67 59.26 63.03 57.84
Table 3: Entropy-threshold sensitivity. Validation accuracy on finance (CFA-L2) and math (MATH-500) as we sweep $\tau\in\{0,0.5,1,1.5,2,2.5,3,4,\infty\}$. Here $\tau{=}0$ branches at every step (equivalent to beam search) and $\tau{=}\infty$ disables branching (self-consistency). Bold indicates settings that outperform both endpoints ($\tau{=}0$ and $\tau{=}\infty$) for the same model and task.
Entropy-threshold sensitivity
We study the sensitivity of EGB to the entropy threshold $\tau$ by sweeping $\tau\in\{0,0.5,1,1.5,2,2.5,3,4,\infty\}$ for six base models on validation sets from two domains (finance: CFA Level II; math: MATH-500). This sweep induces a unified continuum that recovers standard baselines as special cases: $\tau{=}0$ branches at every step (equivalent to beam search), $\tau\in(0,\infty)$ performs selective branching only at high-entropy tokens (EGB), and $\tau{=}\infty$ disables branching entirely (equivalent to self-consistency / best-of-$n$ with voting). Across model families, we observe broad regions of $\tau$ that consistently outperform both endpoints, with accuracy typically varying by only 1–3 percentage points when $\tau$ is perturbed by $\pm 0.5$$1.0$ around a reasonable choice, indicating a practical “safe zone” rather than a sharply tuned optimum. Moreover, the same $\tau$ ranges that perform well on MATH-500 also tend to perform well on CFA-L2 for a given model family, suggesting that $\tau$ behaves primarily as a model-family property rather than a task-specific hyperparameter. While our main configuration uses a single $\tau$ per model family for simplicity and ease of deployment, these sweeps also indicate that lightweight per-task adjustment of $\tau$ can yield additional gains (often on the order of 1–3 points) when tuning on a validation set is feasible.
Thinking-mode models
[…]
Qwen3-1.7B Non-thinking 39.77 50.57 51.70 +1.13 +11.93
Thinking 53.98 54.64 54.00 -0.64 +0.02
Qwen3-4B Non-thinking 54.55 58.27 60.80 +2.53 +6.25
Thinking 69.32 67.01 72.16 +5.15 +2.84
Qwen3-8B Non-thinking 51.70 57.39 61.36 +3.97 +9.66
Thinking 68.75 70.38 72.97 +2.59 +4.22
Table 4: CFA Level II accuracy (%) for Qwen3 models with thinking-mode enabled vs. disabled. We report base (greedy) decoding, Beam Search ($\tau{=}0$), and EGB (selective branching), along with absolute gains of EGB over Beam Search and over the base setting.